OpenStudy (anonymous):

You have a spaceship that is a thin spherical shell. Its mass is 12000 kg and has a radius of 40m. You have one rocket that is located on the surface of the shell. When fired, the shell will begin to rotate. What is the force given by the rocket such that a person standing on the inside of the shell will experience normal gravity at 30 minutes? What is its angular acceleration? How many times did the shell rotate during this time?

OpenStudy (doc.brown):

How do you calculate the moment of inertia for a hollow sphere?

OpenStudy (anonymous):

I= \[\frac{ 2mr ^{2} }{ 3}\]

OpenStudy (doc.brown):

What velocity do you need for an object at 40m to have a radial acceleration of 9.8m/s^2?

OpenStudy (doc.brown):

\[a_R =\frac{v^2}{r}\]\[9.8m/s^2 =\frac{v^2}{40m}\]\[v=20m/s\]

OpenStudy (doc.brown):

\[\omega=\alpha t\]\[20m/s=\alpha 1800s\]\[\alpha=0.011m/s^2\]